# Topological symmetry groups of the Generalized Petersen Graphs

### v1

@@ 4. The case $k^{2}\equiv-1\pmod{n}$
Observe that $\alpha$ interchanges the inner and outer cycles of $P(n,k)$, while $\rho$ induces an order $n$ rotation of both cycles. Thus for any $m<n$, $\rho^{m}$ rotates both cycles while $\rho^{m}\alpha^{\pm 1}$ interchanges them. Furthermore, $\alpha^{2}(u_{i})=\alpha(v_{ki})=u_{k^{2}i}=u_{-i}$ and $\alpha^{2}(v_{i})=v_{-i}$. Thus $\alpha^{2}$ is an involution which turns over both the inner and outer cycles, and hence so is any $\rho^{m}\alpha^{2}$. It follows that every element of the form $\rho^{m}\alpha^{\pm 1}$ has order $4$. Also, since $n$ is not divisible by $4$, $D_{n}=\langle\rho,\alpha^{2}\rangle$ has no elements of order $4$. Thus an element of $B(n,k)$ has order $4$ if and only if it can be expressed as $\rho^{m}\alpha^{\pm 1}$.
@@
Suppose that $n$ is odd. Since $\beta\in B(n,k)$, it follows that $h$ must leave at least one spoke $e_{i}$ setwise invariant. But this implies that $h$ fixes the midpoint of $e_{i}$, which contradicts the above paragraph. Thus no order $4$ element of $B(n,k)$ is positively realizable. ∎
@@
By Theorem 2.1, $D_{n}$ and all of its subgroups are positively realizable for $P(n,k)$. If $H$ is not contained in $D_{n}$, then $H$ must contain an element of order $4$. When $n$ is odd, no order $4$ element of $B(n,k)$ is positively realizable, and hence $H$ is not positively realizable.
Thus we assume that $n$ is even. Since $n$ is not divisible by $4$, we have $n\equiv{2}\pmod{4}$. In order to construct an embedding of $P(n,k)$ which positively realizes $\mathbb{Z}_{4}\leq B(n,k)$, we consider the glide rotation $h$ which rotates a solid torus meridionally by $\frac{\pi}{2}$ while rotating it longitudinally by $\pi$.
Let $U$ be a meridian of the solid torus with $n-2$ evenly spaced vertices labeled $u_{1},\dots,u_{\frac{n}{2}-1},u_{\frac{n}{2}+1},\dots,u_{n-1}$. Let $V=h(U)$ with vertices $v_{k},v_{2k},\dots,v_{\frac{n}{2}-k},v_{\frac{n}{2}}+k,\dots,v_{n-k}$ such that each $v_{j}$ is $h(u_{i})$ for some $i$. Let $D_{U}$ and $D_{V}$ denote disjoint meridional disks bounded by $U$ and $V$ respectively, such that $h(D_{U})=D_{V}$. Let $x$ denote the midpoint of an arc $A$ of $C-(D_{U}\cup D_{V})$. Let $u_{n}$ be a vertex on the arc of $A-\{x\}$ with one endpoint on $D_{U}$, and let $v_{n}$ be a vertex on the arc of $A-\{x\}$ with one endpoint on $D_{V}$. Finally, let $v_{\frac{n}{2}}=h(u_{n})$ and $u_{\frac{n}{2}}=h(v_{n})$. This gives us an embedded set of vertices $W$. Figure 2 illustrates the embedded vertices of $P(10,3)$ except for $v_{5}$ and $u_{5}$, which are in the front of the solid torus. Note that the core is illustrated in green, and $h$ takes the pair of blue arcs on $U$ to the pair of blue arcs on $V$.
@@
Let $G=\langle h\rangle=\mathbb{Z}_{4}$. Then $G$ induces a faithful action $H\leq B(n,k)$ of $W$ such that no pair of adjacent vertices are fixed by a non-trivial element of $H$. The pairs of adjacent vertices $u_{n}$ and $v_{n}$ and $u_{\frac{n}{2}}$ and $v_{\frac{n}{2}}$ are the only ones which are fixed by $h^{2}$. But they are not fixed by any other non-trivial element of $G$. Also these pairs each bound an arc in $C=\mathrm{fix}(h^{2})$ which is disjoint from the other vertices. Hence we can apply Lemma 1.7 to embed the edges of $P(10,3)$ such that the resulting embedded graph is setwise invariant under $G$.
@@
Thus we can apply Theorem 1.5 to conclude that $H^{\prime}$ is either a subgroup of $D_{m}\times D_{m}$ for some odd $m$ or a finite subgroup of $\mathrm{SO}(3)$. However, since $H\leq B(n,k)$ and contains $\mathbb{Z}_{4}$ as a proper subgroup, $H$ has the form $\mathrm{Z}_{r}\rtimes\mathbb{Z}_{4}$. But this is impossible since $H^{\prime}\cong H$. It follows that no subgroup of $B(n,k)$ containing $\mathbb{Z}_{4}$ as a proper subgroup is positively realizable for $P(n,k)$. ∎
@@
Let $\Lambda$ denote the embedding of $P(n,k)$ given at the beginning of Section 3. Let $h$ be an order $2$ rotation of $S^{3}$ interchanging the positively oriented $U$ with the positively oriented $V$ such that it interchanges $u_{i}$ and $v_{ki}$. Now $h(u_{ki})=v_{k^{2}i}=v_{-i}$, and hence $h(v_{i})=u_{-ki}$. Also, let $R$ be a reflection through a sphere containing the circle $V$ which leaves $U$ setwise invariant, fixing $u_{n}$ and its antipodal point on $U$ (which is a vertex if $n$ is even). Thus for every $i$, we have $R(u_{i})=u_{-i}$ and $R(v_{i})=v_{i}$. While both $h$ and $R$ take edges contained in $U$ and $V$ to other such edges, neither $h$ nor $R$ takes edges of the form $\overline{u_{i}v_{i}}$ to other edges. Thus we are instead interested in the orientation reversing isometry $Rh$. Observe that $Rh(u_{i})=R(v_{ki})=v_{ki}$ and $Rh(v_{i})=R(u_{-ki})=u_{ki}$. Thus for each spoke $e_{i}$, we have $Rh(e_{i})=e_{ki}$. It follows that $Rh$ takes $\Lambda$ to itself interchanging $U$ and $V$ and $(Rh)^{2}$ turns over $U$ and $V$, fixing $u_{n}$ and $v_{n}$. In particular, $Rh$ induces $\alpha$ on $\Lambda$.
@@
If $(n,k)$ is non-exceptional then $\mathrm{Aut}(P(n,k))=B(n,k)$, and hence we have the following.
@@
Let $k^{2}\equiv-1\pmod{n}$ and $(n,k)$ be non-exceptional. Then $\mathrm{Aut}(P(n,k))$ and all of its subgroups are realizable for $P(n,k)$. If $n$ is odd then the only groups which are positively realizable for $P(n,k)$ are $D_{n}$ and its subgroups, and if $n$ is even, then the only groups which are positively realizable for $P(n,k)$ are $\mathbb{Z}_{4}$ and $D_{n}$ and its subgroups.
@@ 5. The exceptional case $P(4,1)$
Observe that the only non-trivial automorphisms of $P(4,1)$ which pointwise fix an edge, interchange two pairs of non-adjacent vertices while pointwise fixing two non-adjacent edges.
@@
Figure 3. $\Gamma$ is obtained from the left image by identifying vertices with the same labels and pairs of adjacent branched edges. On the right, we see $\overline{v_{1}v_{2}v_{3}v_{4}}$ is the connected sum of four copies of $J$ and four identical $3_{1}$ knots.
@@
Observe that the rotations of a solid cube leave $\Gamma$ invariant, inducing the automorphisms $\rho=(u_{1},u_{2},u_{3},u_{4})(v_{1},v_{2},v_{3},v_{4})$ and $\sigma=(u_{1}v_{2}v_{4})(u_{2}v_{3}u_{4})$ which generate $S_{4}$. Thus $S_{4}=\langle\rho,\sigma\rangle\leq\mathrm{TSG}_{+}(\Gamma)$. Let $\delta=(u_{1},u_{3})(v_{1},v_{3})$ and note that $\operatorname{Aut}((P(4,1))=S_{4}\times\mathbb{Z}_{2}=\langle\rho,\sigma,\delta\rangle$ has $S_{4}=\langle\rho,\sigma\rangle$ as a maximal subgroup. We show below that $\delta\not\in\mathrm{TSG}_{+}(\Gamma)$.
@@
Let $\Gamma$ be the embedding of $P(4,1)$ whose unfolded projection on the faces of a cube is illustrated on the left in Figure 4. The projection on a single face is illustrated on the right.
Figure 4. The embedding $\Gamma$ of $P(4,1)$ is obtained from the projection on the left by identifying vertices with the same labels.
@@
Observe that no homeomorphism can induce $(u_{1}u_{2}u_{3}u_{4})(v_{1}v_{2}v_{3}v_{4})$, since such a homeomorphism would send knotted edges to unknotted edges. Thus we have $\mathrm{TSG}_{+}(\Gamma)=D_{3}\times\mathbb{Z}_{2}$. Furthermore, the edge $\overline{u_{4}v_{4}}$ is not pointwise fixed by any non-trivial element of $\mathrm{TSG}_{+}(\Gamma)$, and hence by Theorem 1.6 every subgroup of $D_{3}\times\mathbb{Z}_{2}$ is positively realizable. ∎
@@ 6. The exceptional case $P(8,3)$
The generalized Petersen graph $P(8,3)$ is also known as the Möbius-Kantor graph. It follows from [11] that $\mathrm{Aut}(P(8,3))=GL(2,3)\rtimes\mathbb{Z}_{2}=\langle\mu,\beta,\gamma\rangle$, where
$\mu$ $=(u_{1},u_{7},v_{8})(u_{2},v_{7},v_{5})(u_{3},v_{4},u_{5})(u_{6},v_{3},v_{1})$
$\beta$ $=(u_{1},u_{7})(u_{2},u_{6})(u_{3},u_{5})(v_{1},v_{7})(v_{2},v_{6})(v_{3},v_{5})$
$\gamma$ $=(u_{1},u_{2},u_{3},u_{4},u_{5},u_{6},u_{7},u_{8})(v_{1},v_{2},v_{3},v_{4},v_{5},v_{6},v_{7},v_{8}).$
@@
Figure 5. An embedding $\Gamma$ with $\mathrm{TSG}_{+}(\Gamma)=\mathrm{Aut}(P(8,3))$.
@@
Let $\beta=(u_{1},u_{7})(u_{2},u_{6})(u_{3},u_{5})(v_{1},v_{7})(v_{2},v_{6})(v_{3},v_{5})$
$\beta_{1}=(u_{1},u_{2},u_{3},v_{3},v_{8},v_{5},u_{5},u_{6},u_{7},v_{7},v_{4},v_{1})(u_{4},v_{6},u_{8},v_{2}).$
Then $D_{12}=\langle\beta_{1},\beta\rangle$ leaves the set of vertices $U=\{u_{4},v_{6},u_{8},v_{2}\}$ invariant, but $\mathrm{Aut}(P(8,3))$ does not leave $U$ invariant, (as in Figure 5). We create an embedding $\Gamma_{12}$ by adding identical $3_{1}$ knots to the edges of $\Gamma$ that include a vertex in $U$. Then $\langle\beta_{1},\beta\rangle\leq\mathrm{TSG}_{+}(\Gamma_{12})$. However, $D_{12}$ is a maximal subgroup of $GL(2,3)\rtimes\mathbb{Z}_{2}$. Thus $\mathrm{TSG}_{+}(\Gamma_{12})=D_{12}$.
@@
Let $A=\{u_{i},v_{j}\ |\ i\text{ even, }j\text{ odd}\}$ and $B=\{u_{i},v_{j}\ |\ i\text{ odd, }j\text{ even}\}$. Starting with the embedding $\Gamma$ from Figure 5, we add the non-invertible knot $8_{17}$ to each edge between a vertex in $A$ and a vertex in $B$ oriented from $A$ to $B$. This gives us an embedding $\Gamma_{GL}$.
Now $GL(2,3)=\langle\delta_{1},\delta_{2}\rangle$ for
$\delta_{1}$ $=(u_{2},v_{1},u_{8})(u_{3},v_{6},v_{8})(u_{4},u_{6},v_{5})(u_{7},v_{2},v_{4})$
$\delta_{2}$ $=(u_{1},u_{3},v_{4},v_{2},u_{5},u_{7},v_{8},v_{6})(u_{2},u_{4},v_{7},v_{5},u_{6},u_{8},v_{3},v_{1}).$
Observe that $A$ and $B$ are each invariant under $\langle\delta_{1},\delta_{2}\rangle$. Thus $\langle\delta_{1},\delta_{2}\rangle\leq\mathrm{TSG}_{+}(\Gamma_{GL})$. On the other hand, $\gamma=(u_{1},u_{2},u_{3},u_{4},u_{5},u_{6},u_{7},u_{8})(v_{1},v_{2},v_{3},v_{4},v_{5},v_{6},v_{7},v_{8})$ does not leave $A$ and $B$ invariant. Thus $\gamma$ cannot be induced by a homeomorphism of $(S^{3},\Gamma_{GL})$, and hence $\gamma\not\in\mathrm{TSG}_{+}(\Gamma_{GL})$. Since $\langle\delta_{1},\delta_{2}\rangle=GL(2,3)$ is a maximal subgroup of $\mathrm{Aut}(P(8,3))=GL(2,3)\rtimes\mathbb{Z}_{2}$, we have $\mathrm{TSG}_{+}(\Gamma_{GL})=GL(2,3)$. ∎
@@
Figure 6. We embed the edges near $u_{1}$ as on the left. An edge $e_{1}$ is knotted around an edge $e_{2}$ if there is a tangle as on the right.
@@
$\rho_{1}$ $=(u_{2},v_{1},u_{8})(u_{3},v_{6},v_{8})(u_{4},u_{6},v_{5})(u_{7},v_{2},v_{4})$
$\rho_{2}$ $=(u_{1},v_{3},u_{5},v_{7})(u_{2},v_{8},u_{6},v_{4})(u_{3},v_{5},u_{7},v_{1})(u_{4},v_{2},u_{8},v_{6})$
$\rho_{3}$ $=(u_{1},u_{7},v_{8})(u_{2},v_{7},v_{5})(u_{3},v_{4},u_{5})(u_{6},v_{3},v_{1}).$
Now, since $\langle\rho_{1},\rho_{2},\rho_{3}\rangle$ acts transitively on the vertices, we can apply the isometry group $G$ to the neighborhood $N(u_{1})$ to modify a neighborhood around every vertex of $\Gamma^{\prime}$. To check that this is well-defined, we used Sage to determine that the only non-trivial automorphisms in $\langle\rho_{1},\rho_{2},\rho_{3}\rangle$ which fix a vertex are of order $3$. Hence no edge is pointwise fixed by any non-trivial element of $\langle\rho_{1},\rho_{2},\rho_{3}\rangle$. It follows that no automorphism in $\langle\rho_{1},\rho_{2},\rho_{3}\rangle$ changes the order of the edges sharing a vertex. Thus our modification of the neighborhoods around vertices is well-defined, and hence this gives us a well-defined embedding $\Gamma_{SL2}$ of $P(8,3)$ such that the edges around every vertex have a fixed order. By our construction, $SL(2,3)\rtimes\mathbb{Z}_{2}=\langle\rho_{1},\rho_{2},\rho_{3}\rangle\leq\mathrm{TSG}_{+}(\Gamma_{SL2})$.
Observe that $\beta=(u_{1},u_{7})(u_{2},u_{6})(u_{3},u_{5})(v_{1},v_{7})(v_{2},v_{6})(v_{3},v_{5})\in\mathrm{Aut}(P(8,3)=GL(2,3)\rtimes\mathbb{Z}_{2}$ does not preserve the order of edges around $u_{4}$ since $\beta$ fixes $\overline{u_{4}u_{5}}$ and interchanges $u_{3}$ and $u_{5}$. Thus $\mathrm{TSG}_{+}(\Gamma_{SL2})\not=\mathrm{Aut}(P(8,3))$. Since $SL(2,3)\rtimes\mathbb{Z}_{2}$ is a maximal subgroup of $\mathrm{Aut}(P(8,3))$, it follows that $\langle\rho_{1},\rho_{2},\rho_{3}\rangle=\mathrm{TSG}_{+}(\Gamma_{SL2})$. Now since no edge of $P(8,3)$ is pointwise fixed by any non-trivial element of $\langle\rho_{1},\rho_{2},\rho_{3}\rangle$, we can apply Theorem 1.6 to conclude that all of the subgroups of $SL(2,3)\rtimes\mathbb{Z}_{2}$ are also positively realizable. ∎

### v2

@@ 4. The case $k^{2}\equiv-1\pmod{n}$
Since $\alpha$ interchanges the inner and outer cycles of $P(n,k)$ while $\rho$ induces an order $n$ rotation of both cycles, for any $m<n$, $\rho^{m}$ rotates both cycles while $\rho^{m}\alpha^{\pm 1}$ interchanges them. Furthermore, $\alpha^{2}(u_{i})=\alpha(v_{ki})=u_{k^{2}i}=u_{-i}$ and $\alpha^{2}(v_{i})=v_{-i}$. Thus $\alpha^{2}$ is an involution which turns over both the inner and outer cycles, and hence so is any $\rho^{m}\alpha^{2}$. It follows that every element of the form $\rho^{m}\alpha^{\pm 1}$ has order $4$. Also, since $n$ is not divisible by $4$, $D_{n}=\langle\rho,\alpha^{2}\rangle$ has no elements of order $4$. Thus an element of $B(n,k)$ has order $4$ if and only if it can be expressed as $\rho^{m}\alpha^{\pm 1}$.
@@
Suppose that $n$ is odd. Since $\beta\in B(n,k)$, $h$ must leave at least one spoke $e_{i}$ setwise invariant. But this implies that $h$ fixes the midpoint of $e_{i}$, which contradicts the above paragraph. Thus no order $4$ element of $B(n,k)$ is positively realizable. ∎
@@
By Theorem 2.1, $D_{n}$ and all of its subgroups are positively realizable for $P(n,k)$. If $H$ is not contained in $D_{n}$, then $H$ must contain an element of order $4$. When $n$ is odd, no order $4$ element of $B(n,k)$ is positively realizable by Lemma 4.1, and hence $H$ is not positively realizable.
Thus we assume that $n$ is even. Since $n$ is not divisible by $4$, we have $n\equiv{2}\pmod{4}$. In order to construct an embedding of $P(n,k)$ which positively realizes $\mathbb{Z}_{4}\leq B(n,k)$, we consider the glide rotation $h$ which rotates a standardly embedded solid torus meridionally by $\frac{\pi}{2}$ while rotating it longitudinally by $\pi$.
Let $U$ be a meridian of the solid torus with $n-2$ evenly spaced vertices labeled $u_{1},\dots,u_{\frac{n}{2}-1},u_{\frac{n}{2}+1},\dots,u_{n-1}$. Let $V=h(U)$ with vertices $v_{k},v_{2k},\dots,v_{\frac{n}{2}-k},v_{\frac{n}{2}}+k,\dots,v_{n-k}$ such that each $v_{j}$ is $h(u_{i})$ for some $i$. Let $D_{U}$ and $D_{V}$ denote disjoint meridional disks bounded by $U$ and $V$ respectively, such that $h(D_{U})=D_{V}$. Let $C=\mathrm{fix}(h^{2})$, and let $x$ denote the midpoint of an arc $A$ of $C-(D_{U}\cup D_{V})$. Let $u_{n}$ be a vertex on the arc of $A-\{x\}$ with one endpoint on $D_{U}$, and let $v_{n}$ be a vertex on the arc of $A-\{x\}$ with one endpoint on $D_{V}$. Finally, let $v_{\frac{n}{2}}=h(u_{n})$ and $u_{\frac{n}{2}}=h(v_{n})$. This gives us an embedded set of vertices $W$.
For example, Figure 3 illustrates the embedded vertices of $P(10,3)$ except for $v_{5}$ and $u_{5}$, which are in the front of the solid torus. Note that the core is illustrated in green, and $h$ takes the pair of blue arcs on $U$ to the pair of blue arcs on $V$.
@@
Let $G=\langle h\rangle=\mathbb{Z}_{4}$. Then $G$ induces a faithful action $H\leq B(n,k)$ of $W$ such that no pair of adjacent vertices are fixed by a non-trivial element of $H$. The pairs of adjacent vertices $u_{n}$ and $v_{n}$ and $u_{\frac{n}{2}}$ and $v_{\frac{n}{2}}$ are the only ones which are fixed by $h^{2}$. But they are not fixed by any other non-trivial element of $G$. Also these pairs each bound an arc in $C$ which is disjoint from the other vertices. Hence we can apply Lemma 1.8 to embed the edges of $P(n,k)$ such that the resulting embedded graph is setwise invariant under $G$.
@@
Thus we can apply Theorem 1.6 to conclude that $H^{\prime}$ is either a subgroup of $D_{m}\times D_{m}$ for some odd $m$ or a finite subgroup of $\mathrm{SO}(3)$. However, since $H\leq B(n,k)$ and contains $\mathbb{Z}_{4}$ as a proper subgroup, $H$ has the form $\mathbb{Z}_{r}\rtimes\mathbb{Z}_{4}$. But this is impossible since $H^{\prime}\cong H$. It follows that no subgroup of $B(n,k)$ containing $\mathbb{Z}_{4}$ as a proper subgroup is positively realizable for $P(n,k)$. ∎
@@
Let $\Lambda$ denote the embedding of $P(n,k)$ from the beginning of Section 3. Let $h$ be an order $2$ rotation of $S^{3}$ interchanging the positively oriented $U$ with the positively oriented $V$ such that it interchanges $u_{i}$ and $v_{ki}$. Now $h(u_{ki})=v_{k^{2}i}=v_{-i}$, and hence $h(v_{i})=u_{-ki}$. Also, let $R$ be a reflection through a sphere containing the circle $V$ which leaves $U$ setwise invariant, fixing $u_{n}$ and its antipodal point on $U$ (which is a vertex if $n$ is even). Thus for every $i$, we have $R(u_{i})=u_{-i}$ and $R(v_{i})=v_{i}$. While both $h$ and $R$ take edges contained in $U$ and $V$ to other such edges, neither $h$ nor $R$ takes edges of the form $\overline{u_{i}v_{i}}$ to other edges. Thus we are instead interested in the orientation reversing isometry $Rh$. Observe that $Rh(u_{i})=R(v_{ki})=v_{ki}$ and $Rh(v_{i})=R(u_{-ki})=u_{ki}$. Thus for each spoke $e_{i}$, we have $Rh(e_{i})=e_{ki}$. It follows that $Rh$ takes $\Lambda$ to itself interchanging $U$ and $V$ and $(Rh)^{2}$ turns over $U$ and $V$, fixing $u_{n}$ and $v_{n}$. In particular, $Rh$ induces $\alpha$ on $\Lambda$.
@@
If $(n,k)$ is non-exceptional then $\mathrm{Aut}(P(n,k))=B(n,k)$, implying the following.
@@
Let $k^{2}\equiv-1\pmod{n}$ and suppose that $(n,k)$ is non-exceptional. Then $\mathrm{Aut}(P(n,k))$ and all of its subgroups are realizable for $P(n,k)$. If $n$ is odd then the only groups which are positively realizable for $P(n,k)$ are $D_{n}$ and its subgroups, and if $n$ is even, then the only groups which are positively realizable for $P(n,k)$ are $\mathbb{Z}_{4}$ and $D_{n}$ and its subgroups.
@@ 5. The exceptional case $P(4,1)$
The non-trivial automorphisms of $P(4,1)$ which pointwise fix an edge, interchange two pairs of non-adjacent vertices and pointwise fix two non-adjacent edges.
@@
Figure 4. The embedding $\Gamma$ of $P(4,1)$ with $\mathrm{TSG}_{+}(\Gamma)=S_{4}$ is obtained from the left image by identifying vertices with the same labels and pairs of adjacent branched edges. On the right, we see that $\overline{v_{1}v_{2}v_{3}v_{4}}$ is the connected sum of four copies of $J$ and four trefoils.
@@
Observe that the rotations of a solid cube leave $\Gamma$ invariant, inducing the automorphisms $\rho=(u_{1}u_{2}u_{3}u_{4})(v_{1}v_{2}v_{3}v_{4})$ and $\sigma=(u_{1}v_{2}v_{4})(u_{2}v_{3}u_{4})$ which generate $S_{4}$. Thus $S_{4}=\langle\rho,\sigma\rangle\leq\mathrm{TSG}_{+}(\Gamma)$. Let $\delta=(u_{1}u_{3})(v_{1}v_{3})$ and note that $\operatorname{Aut}((P(4,1))=S_{4}\times\mathbb{Z}_{2}=\langle\rho,\sigma,\delta\rangle$ has $S_{4}=\langle\rho,\sigma\rangle$ as a maximal subgroup. We show below that $\delta\not\in\mathrm{TSG}_{+}(\Gamma)$.
@@
Let $\Gamma$ be the embedding of $P(4,1)$ whose unfolded projection on the faces of a cube is illustrated on the left in Figure 5, with the projection on a single face illustrated on the right.
Figure 5. The embedding $\Gamma$ of $P(4,1)$ with $\mathrm{TSG}_{+}(\Gamma)=A_{4}\times\mathbb{Z}_{2}$ is obtained from the projection on the left by identifying vertices with the same labels. On the right, we have the projection on a single face.
@@
No homeomorphism can induce $(u_{1}u_{2}u_{3}u_{4})(v_{1}v_{2}v_{3}v_{4})$, since such a homeomorphism would send knotted edges to unknotted edges. Thus we have $\mathrm{TSG}_{+}(\Gamma)=D_{3}\times\mathbb{Z}_{2}$. Furthermore, the edge $\overline{u_{4}v_{4}}$ is not pointwise fixed by any non-trivial element of $\mathrm{TSG}_{+}(\Gamma)$, and hence by Theorem 1.7 every subgroup of $D_{3}\times\mathbb{Z}_{2}$ is positively realizable. ∎
@@ 6. The exceptional case $P(8,3)$
It follows from [11] that $\mathrm{Aut}(P(8,3))=GL(2,3)\rtimes\mathbb{Z}_{2}=\langle\mu,\beta,\gamma\rangle$, where
$\mu$ $=(u_{1}u_{7}v_{8})(u_{2}v_{7}v_{5})(u_{3}v_{4}u_{5})(u_{6}v_{3}v_{1})$
$\beta$ $=(u_{1}u_{7})(u_{2}u_{6})(u_{3}u_{5})(v_{1}v_{7})(v_{2}v_{6})(v_{3}v_{5})$
$\gamma$ $=(u_{1}u_{2}u_{3}u_{4}u_{5}u_{6}u_{7}u_{8})(v_{1}v_{2}v_{3}v_{4}v_{5}v_{6}v_{7}v_{8}).$
@@
Figure 6. An embedding $\Gamma$ of $P(8,3)$ with $\mathrm{TSG}_{+}(\Gamma)=\mathrm{Aut}(P(8,3))$.
@@
Let $\beta=(u_{1}u_{7})(u_{2}u_{6})(u_{3}u_{5})(v_{1}v_{7})(v_{2}v_{6})(v_{3}v_{5})$
$\beta_{1}=(u_{1}u_{2}u_{3}v_{3}v_{8}v_{5}u_{5}u_{6}u_{7}v_{7}v_{4}v_{1})(u_{4}v_{6}u_{8}v_{2}).$
Then $D_{12}=\langle\beta_{1},\beta\rangle$ leaves the set of vertices $U=\{u_{4},v_{6},u_{8},v_{2}\}$ invariant, but $\mathrm{Aut}(P(8,3))$ does not leave $U$ invariant. We create an embedding $\Gamma_{12}$ by adding identical $3_{1}$ knots to the edges of $\Gamma$ that include a vertex in $U$. Then $\langle\beta_{1},\beta\rangle\leq\mathrm{TSG}_{+}(\Gamma_{12})$. Since $D_{12}$ is a maximal subgroup of $GL(2,3)\rtimes\mathbb{Z}_{2}$, $\mathrm{TSG}_{+}(\Gamma_{12})=D_{12}$.
@@
Let $A=\{u_{i},v_{j}\ |\ i\text{ even, }j\text{ odd}\}$ and $B=\{u_{i},v_{j}\ |\ i\text{ odd, }j\text{ even}\}$. Starting with the embedding $\Gamma$ from Figure 6, we add the non-invertible knot $8_{17}$ to each edge between a vertex in $A$ and a vertex in $B$ oriented from $A$ to $B$. This gives us an embedding $\Gamma_{GL}$. Now $GL(2,3)=\langle\delta_{1},\delta_{2}\rangle$ for
$\delta_{1}$ $=(u_{2}v_{1}u_{8})(u_{3}v_{6}v_{8})(u_{4}u_{6}v_{5})(u_{7}v_{2}v_{4})$
$\delta_{2}$ $=(u_{1}u_{3}v_{4}v_{2}u_{5}u_{7}v_{8}v_{6})(u_{2}u_{4}v_{7}v_{5}u_{6}u_{8}v_{3}v_{1}).$
Observe that $A$ and $B$ are each invariant under $\langle\delta_{1},\delta_{2}\rangle$. Thus $\langle\delta_{1},\delta_{2}\rangle\leq\mathrm{TSG}_{+}(\Gamma_{GL})$. On the other hand, $\gamma=(u_{1}u_{2}u_{3}u_{4}u_{5}u_{6}u_{7}u_{8})(v_{1}v_{2}v_{3}v_{4}v_{5}v_{6}v_{7}v_{8})$ does not leave $A$ and $B$ invariant. Thus $\gamma$ cannot be induced by a homeomorphism of $(S^{3},\Gamma_{GL})$, and hence $\gamma\not\in\mathrm{TSG}_{+}(\Gamma_{GL})$. Since $\langle\delta_{1},\delta_{2}\rangle=GL(2,3)$ is a maximal subgroup of $\mathrm{Aut}(P(8,3))=GL(2,3)\rtimes\mathbb{Z}_{2}$, we have $\mathrm{TSG}_{+}(\Gamma_{GL})=GL(2,3)$. ∎
@@
Figure 7. We embed the edges near $u_{1}$ as on the left. We say $e_{1}$ is knotted around $e_{2}$ if there is a tangle as on the right.
@@
$\rho_{1}$ $=(u_{2}v_{1}u_{8})(u_{3}v_{6}v_{8})(u_{4}u_{6}v_{5})(u_{7}v_{2}v_{4})$
$\rho_{2}$ $=(u_{1}v_{3}u_{5}v_{7})(u_{2}v_{8}u_{6}v_{4})(u_{3}v_{5}u_{7}v_{1})(u_{4}v_{2}u_{8}v_{6})$
$\rho_{3}$ $=(u_{1}u_{7}v_{8})(u_{2}v_{7}v_{5})(u_{3}v_{4}u_{5})(u_{6}v_{3}v_{1}).$
Since $\langle\rho_{1},\rho_{2},\rho_{3}\rangle$ acts transitively on the vertices, we can apply the isometry group $G$ to the neighborhood $N(u_{1})$ to modify a neighborhood around every vertex of $\Gamma^{\prime}$. To check that this is well-defined, we used Sage to determine that the only non-trivial automorphisms in $\langle\rho_{1},\rho_{2},\rho_{3}\rangle$ which fix a vertex are of order $3$. Hence no edge is pointwise fixed by any non-trivial element of $\langle\rho_{1},\rho_{2},\rho_{3}\rangle$. It follows that no automorphism in $\langle\rho_{1},\rho_{2},\rho_{3}\rangle$ changes the order of the edges sharing a vertex. Thus our modification of the neighborhoods around vertices is well-defined, and hence we have a well-defined embedding $\Gamma_{SL2}$ of $P(8,3)$ such that the edges around every vertex have an order. Hence $SL(2,3)\rtimes\mathbb{Z}_{2}=\langle\rho_{1},\rho_{2},\rho_{3}\rangle\leq\mathrm{TSG}_{+}(\Gamma_{SL2})$.
Now $\beta=(u_{1}u_{7})(u_{2}u_{6})(u_{3}u_{5})(v_{1}v_{7})(v_{2}v_{6})(v_{3}v_{5})\in\mathrm{Aut}(P(8,3)=GL(2,3)\rtimes\mathbb{Z}_{2}$ does not preserve the order of edges around $u_{4}$ since $\beta$ fixes $\overline{u_{4}u_{5}}$ and interchanges $u_{3}$ and $u_{5}$. Thus $\mathrm{TSG}_{+}(\Gamma_{SL2})\not=\mathrm{Aut}(P(8,3))$. Since $SL(2,3)\rtimes\mathbb{Z}_{2}$ is a maximal subgroup of $\mathrm{Aut}(P(8,3))$, it follows that $\langle\rho_{1},\rho_{2},\rho_{3}\rangle=\mathrm{TSG}_{+}(\Gamma_{SL2})$. Now since no edge of $P(8,3)$ is pointwise fixed by any non-trivial element of $\langle\rho_{1},\rho_{2},\rho_{3}\rangle$, we can apply Theorem 1.7 to conclude that all of the subgroups of $SL(2,3)\rtimes\mathbb{Z}_{2}$ are also positively realizable. ∎